OpenStudy (anonymous):

derivative of sqrt(1 - x^2)

OpenStudy (anonymous):

\[\sqrt(1 - x^2) \]

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{1-x^2}=\frac{1}{2\sqrt{1-x^2}}\times -2x=\frac{-x}{\sqrt{1-x^2}}\]

OpenStudy (anonymous):

in general the derivative of \[\sqrt{f(x)}\] is \[\frac{f'(x)}{2\sqrt{f(x)}}\] by the chain rule

OpenStudy (anonymous):

cool... that is what I am doing is the chain rule